### v2

@@ 4.4 Analysis of the Proposed Shortcut Connection
Next, we measure the L2 distance (similarity) between each token’s preceding-layer and current-layer representations. Figure 11 (b) illustrates that, as training advances, the L2 similarity initially decreases with network depth, then increases, and ultimately reaches its maximum value in the final layer. Since the gating network is used to classify the representations, this similarity may lead to the repeated selection of the same experts, as evidenced by the correlation between the results in Figure 11 (a) and (b).
@@
Notably, we observe this experimental architecture only achieves model quality on par with the standard top-1 MoE, despite incurring the same computational cost as the top-2 MoE. To address this, we impose a constraint on the MoE module to ensure that it selects a different expert for the current layer than the one chosen for the preceding layer. We refer to this enhanced experimental architecture as DoubleGating MoE (DGMoE), which achieves model quality comparable to standard top-2 MoE by imposing the constraint against selecting the same expert repeatedly.
With this constraint, we observe gating score behaviors are similar to those of the standard top-2 MoE Riquelme et al. (2021), as illustrated in Figure 11 (c) and (d). Specifically, in (c), the gating scores associated with representations from preceding layers grow quickly initially, converging thereafter, and achieving their maximum in the final MoE sub-layer. By comparison, in (d), the gating scores of current layer across all MoE sub-layers are around $0.2$ (±$\pm 0.04$), lower than the $0.2$ to $0.8$ range for that of preceding layer.
In summary, the experiments indicate that the combined results of two representations, produced by the same expert, are approximately equivalent to those obtained by using a single expert to process only the current-layer representations. This equivalence may arise from the similarity between the preceding-layer and current-layer representations of each token. Furthermore, enforcing two distinct experts can enable the DGMoE architecture to achieve behaviors and model quality similar to those of a standard top-2 MoE.
Analogously, our ScMoE architecture has already allocated distinct experts for the two representations of each token, thereby maintaining a similar behavior to the shared-expert MoE architecture and ensuring comparable model quality. Given the differences between standard MoE and shared-expert MoE, we have observed that DGMoE achieves accuracy levels closer to those of the standard top-2 MoE, while ScMoE aligns more closely with shared-expert MoE, as demonstrated in the results presented in Appendix A.2.
Additionally, our ScMoE model may leverage the differences between the representations of preceding layers and those of the current layer, while still preserving significant similarities. This distinction potentially provides additional information that enhances the training of MoE models, especially considering that MoE models typically require larger training datasets compared to dense MLP models Artetxe et al. (2022); Fedus et al. (2022); Xue et al. (2022).
@@ 5 Conclusion
The inherent dependencies of communication in distributed MoE models hinder parallel optimization techniques to improve execution efficiency. To overcome this, we propose a novel shortcut-connected MoE (ScMoE) architecture, and develop a communication overlapping parallel strategy. Through comprehensive empirical evaluation and theoretical analysis, our approaches demonstrate better execution efficiency while maintaining, or exceeding, the model quality of existing methods. Moreover, we further implement an expert offloading strategy upon our ScMoE to enable memory-limited inference and optimize latency by overlapping expert migration. Additionally, we provide insightful analysis and discussion on our proposed MoE architecture.
@@ A.1 Theoretical Analysis
In this section, we delve deeper into the understanding of our proposed shortcut-connected MoE (ScMoE) architecture, presenting a theoretical foundation focused on the propagation of gradients to guarantee consistent training and preserve model quality. Our analysis is confined to our ScMoE architecture as depicted in Figure 4(b); however, the same principles and derivations can be easily extended to other shortcut-connected MoE architectures. Building upon Equations 7 to 10, we can derive
@@
$+\sum_{i=1}^{N}G(\mathcal{H}^{MH}_{l})_{i}E_{i}(\mathcal{H}^{MH}_{l})\Big{)}$
$\mathcal{H}^{MH}_{l}$ $=\mathcal{H}_{l-1}+\text{MultiHead}^{(l)}(\mathcal{H}_{l-1})$ (15)
It is observable that Equations 14 and 15 share an identical structural expression. Consequently, we consider each pair of Block-MoE and Block-MLP layers as a single entity, and every sub-layer, denoted as $\mathcal{F}$, with its corresponding parameters $\mathcal{W}_{l}$, conforms to the equation
$x_{l+1}=x_{l}+\mathcal{F}_{\mathcal{W}_{l}}(x_{l})$ (16)
@@
$\frac{\partial{\mathcal{E}}}{\partial{x_{l}}}$ $=\frac{\partial{\mathcal{E}}}{\partial{x_{L}}}\frac{\partial{x_{L}}}{\partial{x_{l}}}=\frac{\partial{\mathcal{E}}}{\partial{x_{L}}}\Big{(}1+\frac{\partial{}}{\partial{x_{l}}}\sum_{i=1}^{L-1}\mathcal{F}_{\mathcal{W}_{i}}(x_{i})\Big{)}$ (18)
@@ A.2 Analysis of the DoubleGating MoE (DGMoE)
A.3 Analysis of Shared Expert Gate
In contrast to the gate-selected expert, the shared expert is fixed to process all representations without computing a gating score through the gating network. Therefore, common practices for shared-expert MoE employ a shared-expert gate to generate the coefficient for the shared expert’s output Team (2024); Dai et al. (2024). Specifically, the shared-expert gate is a linear layer that uses the input representation of the MoE module as its input to generate a coefficient similar to a gating score. Building upon Equation 6, the output of the shared expert can be expressed as
$SE(x)=SEGate(x)\cdot\text{MLP}(x)$ (20)
where $SEGate(\cdot)$ denotes the shared-expert gate and $\text{MLP}(\cdot)$ represents the essence of the shared expert.
@@ A.3 Analysis of Shared Expert Gate
As demonstrated in Table 5, the incorporation of the shared-expert gate significantly enhances model performance in our experimental vision tasks. In the absence of the shared-expert gate, the quality of MoE architectures with shared experts declines from that of a standard top-2 MoE to that of a standard top-1 MoE, despite maintaining the same computational volume as the standard top-2 MoE.
@@
Table 5: Comparison of top-1 accuracy on the ImageNet-1K test set for SwinV2-MoE-S models, with and without the shared-expert gate (denoted as “SE-Gate”).
Model With SE-Gate Without SE-Gate
@@
[…]
Our ScMoE 79.38% 80.56%
Table 7: Comparison of zero-shot perplexity on WikiText-103 for our pre-trained GPT2-MoE-Small and GPT2-MoE-Medium models with various architectures.
@@
[…]
Our ScMoE 29.10 17.62
A.4 Evaluation Across Different Model Sizes
@@
A.5 Share MoE Across Multiple Layers via Shortcut Connections
From a certain point of view, our shortcut-connected MoE architectures can be conceptualized as the sharing of one MoE module across multiple transformer layers. Parameter sharing across different layers has been validated as a method to enhance parameter efficiency and improve model quality, as evidenced in existing research Lan et al. (2019); Dehghani et al. (2018); Xue et al. (2022); Huang et al. (2017).
@@ A.5 Share MoE Across Multiple Layers via Shortcut Connections
Table 8: Hyperparameters for GPT-MoE models.
Parameter GPT2-MoE-Small GPT2-MoE-Medium GPT3-MoE-XL
Num. layers 12 24 24
Embedding dim 768 1024 2048
Num. attention heads 12 16 32
Num. experts per layer 8 8 8
Num. parameters 323M 1.1B 4.1B
Context/sequence length 1K 2K 2K
Training tokens 180B 180B 18B
Warmup tokens 100M 100M 10M
Batch size 256 64 32
[…]
LR-Scheduler Inverse_sqrt Inverse_sqrt Inverse_sqrt
Capacity factor 2.00 2.00 2.00
MoE loss coefficient 0.01 0.01 0.01
A.6 Experimental Details
@@
Table 8 summarizes the hyperparameters for training the GPT-MoE models including GPT2-MoE-Small, GPT2-MoE-Medium and GPT3-MoE-XL. We implement our models based on the Fairseq Ott et al. (2019) and Tutel Hwang et al. (2022) frameworks. And we use the Openwebtext dataset Gokaslan & Cohen (2019) for pre-training, with the GPT-2 BPE tokenizer Radford et al. (2019). In addition to evaluating the training curve, we also conduct a zero-shot evaluation on WikiText-103 Merity et al. (2016) dataset. Additionally, we maintained uniform hyperparameters across all models to ensure comparability, and we recognize that optimal parameter tuning may further enhance model performance, a subject for subsequent research.
Table 9 summarizes the hyperparameters for training the Swin-MoE models including SwinV2-MoE-S and SwinV2-MoE-B. We develop our experiments based on the models and MoE framework which are provided by Tutel Hwang et al. (2022), and also follow its hyperparameter configurations. The number of gate-selected experts per MoE module corresponds to the number of GPUs. Specifically, the experiments related to overhead and acceleration analysis in a 2-node (16×A800-NVLink) scenario utilize 16 experts per MoE module, while other cases use 8 experts. Then, we pre-train the models on ImageNet-1K image classification dataset. To maintain the comparability of our experiments, we limit our modifications solely to the MoE architectures and keep the hyperparameters and random seeds consistent. In addition, the experimental results related to efficiency are the averages of multiple samples over different periods.

### v3

@@ 5.1.1 Analysis of Gating Behaviors
Next, we measure the L2 distance (similarity) between each token’s preceding-layer and current-layer representations. Figure 9(b) illustrates that, as training advances, the L2 distance initially decreases with network depth, then increases, and ultimately reaches its maximum value in the final layer. Since the gating network is used to classify the representations, this similarity may lead to the repeated selection of the same experts, as evidenced by the correlation between the results in Figure 9(a) and 9(b).
@@
Furthermore, we trained an experimental MoE model incorporating specialized MoE modules that select the top-1 expert twice for processing each input token’s current-layer and preceding-layer representations. We observe that this experimental architecture achieves a model quality equivalent to the standard top-1 MoE, despite incurring the same computational cost as the top-2 MoE. Interestingly, this architecture can achieve model quality comparable to the standard top-2 MoE by imposing a constraint on the MoE module that ensures the selection of a different expert for the current layer than for the preceding layer. We refer to this enhanced experimental architecture as DoubleGating MoE (DGMoE), with further details provided in Appendix A.2. With this constraint, we observe gating score behaviors are similar to those of the standard top-2 MoE (Riquelme et al., 2021), as illustrated in Figure 9(c) and 9(d).
5.1.2 Analysis of Similarity in Representations
Based on the observations mentioned above, we believe that the similarity between each token’s preceding-layer and current-layer representations is crucial to understanding these outcomes and validating the effectiveness of our proposed ScMoE model. Assuming that the representations of the preceding layer and the current layer are identical, utilizing the same expert to process these two representations is equivalent to employing a single expert to process only the current-layer representations, thereby resulting in model quality comparable to that of the standard top-1 MoE. On the other hand, assigning distinct experts to the two representations of each token is equivalent to activating two experts to process the current-layer representations, thereby achieving model quality similar to that of the standard top-2.
Refer to caption (a) Swin-MoE-Small Refer to caption (b) GPT2-MoE-Medium
Figure 10: Analysis of cosine similarity in intermediate representations. The representations include the input to the first layer (denoted as ’In’) and the outputs of the Attention (e.g., ’1A’) and MLP/MoE (e.g., ’1M’) sublayers within each Transformer block.
Moreover, we analyze the similarities in the intermediate representations of the Swin-MoE-Small and GPT2-MoE-Medium models, which use the standard top-2 MoE, as illustrated in Figure 10(a) and 10(b). It is evident that the representations from adjacent Transformer blocks exhibit a cosine similarity close to 1.0, highlighting their high degree of similarity. Consequently, our proposed ScMoE architecture assigns distinct experts to the two representations of each token (a shared expert for current-layer representations and routed experts for preceding-layer representations), thereby preserving behavior akin to the standard top-2 and shared-expert MoE architectures and ensuring comparable model quality. Similar observations in LLaMA2-MoE and OLMoE (Muennighoff et al., 2024), as demonstrated in Appendix A.9, further confirm the generalizability of our ScMoE to other models.
In addition, we provide a theoretical analysis of our proposed ScMoE architecture in Appendix A.1, elucidating the propagation of gradients to guarantee consistent training and preserve model quality.
5.2 Configuration of ScMoE Architecture
Coefficient Gating Network. In contrast to the gate-routed expert, the shared expert is fixed to process all representations without computing a gating score through the gating network. Therefore, some work (Qwen, 2024; Rajbhandari et al., 2022) employs a coefficient gating network $CG$ to generate the coefficient for combining the outputs of gate-routed and shared experts. Specifically, the coefficient gating network is a linear layer that uses the MoE module’s input representation as its input to generate the coefficient.
We conduct experiments on ScMoE using three distinct methods for combining the outputs of gate-routed experts and shared experts: (1) Direct Add, (2) $CG$-1, and (3) $CG$-2. The Direct Add method, indicated by Equation 6, involves directly summing the outputs from both the shared expert and the gate-routed expert. For each input token $x\in\mathbb{R}^{n}$, the MoE outputs using $CG$-1 and $CG$-2, which generate coefficients for the combination, can be expressed as follows:
$CG$-1:
$\text{coef}=Sigmoid(W_{\text{CG}}\cdot x),\quad{W_{\text{CG}}}\in\mathbb{R}^{1\times n},$ (14)
$MoE(x)=\text{coef}\cdot SE(x)+\sum_{i=1}^{N}G(x)_{i}E_{i}(x),$ (15)
$CG$-2:
$\text{coef}=softmax(W_{\text{CG}}\cdot x),\quad{W_{\text{CG}}}\in\mathbb{R}^{2\times n},$ (16)
$MoE(x)=\text{coef}[0]\cdot SE(x)+\text{coef}[1]\cdot\sum_{i=1}^{N}G(x)_{i}E_{i}(x).$ (17)
As shown in Table 3, the configuration of $CG$-1 achieves the lowest final validation loss among the three combination methods, all of which are set to Pos-2. Moreover, ScMoE models with three configurations consistently outperform both the standard top-2 and the shared-expert MoE (configured with $CG$-2 according to (Rajbhandari et al., 2022)).
Shortcut-connected Position. While the Pos-1 configuration can maximize the potential overlap duration when MoE is placed in every block, using different configurations of shortcut-connected positions (Pos-1, Pos-2 or Pos-3) when placing MoE in every second block will result in variations in overlap duration and model quality. Therefore, we conduct experiments to identify its optimal configuration.
As shown in Table 3, ScMoE (Pos-2) achieves the lowest final validation loss among the configurations tested, all of which utilize the $CG$-1 setup. This outcome mirrors the findings from the vision experiments detailed in Table 1, where Pos-2 also delivers the highest accuracy.
As illustrated in Figure 10(b), the input and intermediate representations of the first layer differ significantly from those of subsequent layers, with the MLP/MoE altering the representations more substantially than Attention. Therefore, we explore modifying the first MoE module to utilize Pos-1 while the remaining modules employ Pos-2, a configuration referred to as Pos-2 (L0 Pos-1). This setup results in a slightly higher loss compared to Pos-2 alone. Observations of varying shortcut-connected positions reveal that the superior performance of Pos-2 suggests the model quality of ScMoE is not necessarily improved by connecting more similar or dissimilar intermediate representations.
Consequently, we select the ScMoE configuration of $CG$-1 and Pos-2 for the experimental GPT2-MoE model, with its evaluations presented in Table 2.
Table 3: Comparison of the final validation loss of GPT-2 MoE pre-training across various MoE methods and configurations.
MoE Method Configuration Final Validation loss ($\downarrow$)
Standard top-2 MoE - 3.270405
Shared-Expert MoE - 3.240592
Our ScMoE (Pos-2) Direct Add 3.236811
$CG$-1 3.224763
$CG$-2 3.232943
Our ScMoE ($CG$-1) Pos-1 3.237530
Pos-2 3.224763
Pos-3 3.241349
Pos-2 (L0 Pos-1) 3.225626
5.3 Optimization for Memory-Limited Inference
Existing studies (Hwang et al., 2024; Yi et al., 2023) offload expert parameters to CPU memory in memory-limited inference scenarios where GPU cannot store the full MoE model. These studies utilize information in preceding layers to predict expert selection for the current MoE layer, enabling early expert migration from CPU to GPU and overlapping it with model computation. In contrast to existing speculative expert migration methods, we implement an expert offloading strategy with overlapping determinate migration, built upon our ScMoE that inherently advances expert selection to the preceding layer. The experimental results demonstrate that our expert offloading strategy reduces peak GPU memory usage by up to 60% and decreases expert migration costs by up to 75% through overlapping with computation. More details are shown in Appendix A.3.
@@ 6 Conclusion
The inherent dependency between communication and computation in conventional distributed MoE models hinders parallel optimization techniques to improve execution efficiency. To address this, we propose a shortcut-connected MoE (ScMoE) architecture, and develop a communication overlapping parallel strategy. Through empirical evaluation and theoretical analysis, our approaches demonstrate better execution efficiency while maintaining or exceeding the model quality of existing methods. In addition, we provide an insightful analysis and discussion of ScMoE.
Acknowledgements
We thank the anonymous reviewers for their valuable comments. This work was supported in part by the Guangdong Provincial Project (No. 2023QN10X252), the Guangdong Basic and Applied Basic Research Foundation (No. 2023A1515110353), and GDIC. This research was conducted on the High-Performance Computing Platform of HKUST(GZ).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Theoretical Analysis
In this section, we delve deeper into the understanding of our proposed shortcut-connected MoE (ScMoE) architecture, presenting a theoretical foundation focused on the propagation of gradients to guarantee consistent training and preserve model quality. Our analysis is confined to the ScMoE (Pos-2) architecture as depicted in Figure 4(b); however, the same principles and derivations can be easily extended to other shortcut-connected MoE architectures. Building upon Equations 7 to 10, we can derive
@@
$+\sum_{i=1}^{N}G(\mathcal{H}^{MH}_{l})_{i}E_{i}(\mathcal{H}^{MH}_{l})\Big{)},$
$\mathcal{H}^{MH}_{l}$ $=\mathcal{H}_{l-1}+\text{MultiHead}^{(l)}(\mathcal{H}_{l-1}).$ (19)
It is observable that Equations 18 and 19 share an identical structural expression. Consequently, we consider each pair of Block-MoE and Block-MLP layers as a single entity, and every sub-layer, denoted as $\mathcal{F}$, with its corresponding parameters $\mathcal{W}_{l}$, conforms to the equation
$x_{l+1}=x_{l}+\mathcal{F}_{\mathcal{W}_{l}}(x_{l}).$ (20)
@@
$\frac{\partial{\mathcal{E}}}{\partial{x_{l}}}$ $=\frac{\partial{\mathcal{E}}}{\partial{x_{L}}}\frac{\partial{x_{L}}}{\partial{x_{l}}}=\frac{\partial{\mathcal{E}}}{\partial{x_{L}}}\Big{(}1+\frac{\partial{}}{\partial{x_{l}}}\sum_{i=1}^{L-1}\mathcal{F}_{\mathcal{W}_{i}}(x_{i})\Big{)}.$ (22)
@@ A.2 Analysis of the DoubleGating MoE (DGMoE)
A.3 Shortcut-connected MoE for Optimizing Memory-Limited Inference
While MoE effectively enhances LLMs in terms of model quality, it faces significant deployment challenges during on-device inference due to high memory demand. A common approach is to offload expert parameters to CPU memory (Shen et al., 2022) in scenarios where GPU memory is insufficient to store the entire MoE model. Moreover, decoder-only models use an autoregressive process for natural language generation (NLG) inference tasks, allowing for per-token processing of MoE. Specifically, only the two activated experts (top-2 gating) for each token need to be transferred from CPU to GPU memory for computation, thereby reducing peak GPU memory usage.
Since the migration of activated expert parameters from CPU to GPU, which occurs after expert selection, blocks expert computation until the transfer is complete, existing studies (Hwang et al., 2024; Yi et al., 2023; Du et al., 2024) have explored prefetching the experts. For instance, Pre-gated MoE (Hwang et al., 2024) uses information from preceding layers to predict expert selection, allowing for preloading of expert parameters into GPU memory, as shown in Figure 12 (a). This method enables overlapping the expert migration duration with the computation of preceding modules. Moreover, speculative expert migration methods adjust only the expert selection process, while expert computation continues along the same data flow of representations as in standard MoE.
However, speculative expert migrations can suffer from estimation inaccuracies, as they deviate from the original logic of pre-trained models, potentially reducing inference accuracy. In contrast, our proposed ScMoE architecture utilizes the gate-routed expert to compute the preceding-layer representations, inherently facilitating early expert migration well before the expert computation in the current layer. This allows us to implement an expert offloading strategy with overlapping determinate migration, maintaining the pre-trained logic.
Additionally, existing expert migration methods cannot be adapted to overlap communication in expert parallelism. This is because they do not decouple dependencies in the data flow of expert processing representations, and therefore cannot adjust the All-to-All communication of these representations
Figure 12: Illustrations of various expert migration methods to improve the efficiency of expert offloading: (a) speculative expert migration, exemplified by Pre-gated MoE (Hwang et al., 2024), and (b) our ScMoE’s determinate expert migration. The red dashed line indicates expert selection and the transfer of expert parameters from CPU memory to GPU memory, while the black or red solid lines represent the data flow of representations processed by the Attention, MLP, and expert modules.
A.3.1 Expert Offloading Strategy
We implement an expert offloading strategy that keeps non-expert and shared expert modules in GPU memory while offloading other gate-routed experts to CPU memory. After the Attention module in the preceding layer generates intermediate representations, the gate determines expert selection and issues asynchronous migration of the activated expert, as illustrated in Figure 12(b). This approach allows expert migration to overlap with the computation duration. Importantly, expert selection in our method adheres to the logic of the pre-trained ScMoE model, without speculation.
Figure 13: Peak GPU memory usage (a) and MoE block latency (b) for various memory-limited inference methods applied to the GPT2-MoE-Medium (8 experts per MoE module) and GPT3-MoE-XL models using ScMoE. “GPU-only” indicates that the entire model is stored in GPU memory. “Offload” refers to our strategy of offloading expert parameters to CPU with blocking expert migration. “Offload-Async” denotes the use of asynchronous expert migration to overlap its duration.
A.3.2 Evaluation
We evaluate our proposed expert offloading strategy on models with our ScMoE (Pos-2) architecture, using a platform with a single A30-PCIe GPU. As demonstrated in Figure 13(a), our expert offloading strategy reduces peak GPU memory usage by 50% for the GPT2-MoE-Medium model and by 60% for the GPT3-MoE-XL model when deployed in the inference scenario using a single A30-PCIe GPU. Furthermore, it is anticipated that models with more gate-routed experts in each MoE module will experience a larger percentage reduction in GPU memory usage.
Since the offloaded expert parameters must be loaded into the GPU memory for expert computation, the blocking execution of this expert migration results in significant overhead. As shown in Figure 13(b), the blocking expert migration introduces an additional overhead of 80% in GPT2-MoE-Medium and 240% in GPT3-MoE-XL, substantially increasing the MoE block latency. To mitigate this issue, our strategy of asynchronously executing the determinate expert migration effectively reduces the additional costs by 75% in GPT2-MoE-Medium and 25% in GPT3-MoE-XL.
Furthermore, it is evident that expanding the model size from Medium to XL significantly raises the cost proportion related to expert migration. This is because the per-token decoding process during inference is memory-bound (Patel et al., 2024; Wu et al., 2024). The larger model size leads to a proportional increase in the duration of memory transfer, without a corresponding increase in computation time.
Table 4: Comparison of validation perplexity and end-to-end speedup analysis of train and inference (one iteration) for our pre-trained GPT3-MoE-XL (Brown et al., 2020) models with various architectures in 8$\times$×A800-NVLink scenario, using standard MoE with top-2 gating as the baseline. “ScMoE-2” refers to the activation of one shared expert and two gate-routed experts.
Model Validation Train Inference
[…]
Standard top-2 17.52 1 1
Our ScMoE 16.46 1.12$\times$ 1.18$\times$
Standard top-3 17.26 0.94$\times$ 0.92$\times$
Our ScMoE-2 16.27 1.05$\times$ 1.08$\times$
A.4 Analysis of More Activated Experts
As increasing the number of activated experts within standard MoE is correlated with enhancements in model quality, we implement this augmentation in our ScMoE by increasing the count of gate-routed experts that process the preceding-layer representations, while maintaining the process of current-layer representations. To investigate the benefits of more activated experts, we implement the ScMoE-2, which employs top-2 experts for the preceding layer and one shared expert for the current layer.
A.5 Coefficient Gating Network in Vision Task
@@
As shown in Table 5, the incorporation of the coefficient gating network significantly enhances model performance in our experimental vision tasks. In the absence of the coefficient gating network, the quality of MoE architectures with shared experts declines from that of a standard top-2 MoE to that of a standard top-1 MoE, despite maintaining the same computational volume as the standard top-2 MoE.
@@
Table 5: Comparison of top-1 accuracy on the ImageNet-1K test set for SwinV2-MoE-S models, using Direct Add and $CG$-1.
Model $CG$-1 Direct Add
Shared-Expert MoE 79.53% 79.02%
@@
Our DGMoE (Pos-2) 79.35% 80.51%
Our ScMoE (Pos-2) 79.38% 80.56%
Table 7: Comparison of zero-shot perplexity on WikiText-103 for our pre-trained GPT2-MoE-Small and GPT2-MoE-Medium (8 experts per MoE module) models with various architectures.
@@
Our DGMoE (Pos-2) 31.52 18.91
Our ScMoE (Pos-2) 29.10 17.62
A.6 Evaluation Across Different Model Sizes
@@
A.7 Share MoE Across Multiple Layers via Shortcut Connections
From a certain point of view, our shortcut-connected MoE architectures can be conceptualized as the sharing of one MoE module across multiple transformer layers. Parameter sharing across different layers has been validated as a method to enhance parameter efficiency and improve model quality, as evidenced in existing research (Lan et al., 2019; Dehghani et al., 2018; Xue et al., 2022; Huang et al., 2017).
@@ A.7 Share MoE Across Multiple Layers via Shortcut Connections
Table 8: Hyperparameters for GPT-MoE and LLaMA2-MoE models.
Parameter GPT2-MoE-Small GPT2-MoE-Medium GPT3-MoE-XL LLaMA2-MoE
Num. layers 12 24 24 24
Embedding dim 768 1024 2048 2048
Num. attention heads 12 16 32 16
Num. KV heads 12 16 32 4
Num. experts per layer 8 16 8 8
MoE frequency 1/2 1/2 1/2 1
Num. parameters 323M 1.7B 4.1B 6.7B
Context/sequence length 1K 2K 2K 2K
Capacity factor 2.00 2.00 2.00 2.00
MoE loss coefficient 0.01 0.01 0.01 0.01
@@
A.8 Experimental Details
Hardware Configurations. To assess the effectiveness of our proposed overlapping strategy for enhancing expert parallelism, we conducted experiments on three hardware configurations: 8$\times$×A30-PCIe, 8$\times$×A800-NVLink and 16$\times$×A800-NVLink (across 2 nodes). These configurations cover scenarios with both high and low communication-to-computation ratios. Additionally, we evaluate our proposed expert offloading strategy on a configuration with a single A30-PCIe GPU.
Experiments on Vision Model. To evaluate the efficacy of our MoE architectures on vision tasks, we conduct experiments on SwinV2-MoE model, which is a state-of-the-art vision transformer model built upon the Tutel MoE framework (Hwang et al., 2023; Liu et al., 2021). Specifically, we pre-train the SwinV2-MoE models with various MoE architectures on ImageNet-1K image classification dataset, and subsequently evaluate their accuracy on the corresponding test set. It is noteworthy that the integration of the MoE module within SwinV2 is confined to stages 3 and 4, with our architectural enhancements being selectively applied to the MoE modules in stage 3—the deepest submodel. Given our hardware constraints, we configure each MoE module with 8 experts, assigning one expert per GPU device. Table 9 summarizes the hyperparameters for training the Swin-MoE models including SwinV2-MoE-S and SwinV2-MoE-B. Specifically, the experiments related to overhead and acceleration analysis in a 2-node (16×A800-NVLink) scenario utilize 16 experts per MoE module, while other cases use 8 experts. To maintain the comparability of our experiments, we limit our modifications solely to the MoE architectures and keep the hyperparameters and random seeds consistent. In addition, the experimental results related to efficiency are the averages of multiple samples over different periods.
Experiments on Language Model. For natural language generation (NLG) tasks, we utilize the standard implementations of GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020) and LLaMA-2 (Touvron et al., 2023) from Fairseq (Ott et al., 2019), augmented with Tutel MoE to construct GPT2-MoE, GPT3-MoE and LLaMA2-MoE models. Specifically, we implement GPT2-MoE and GPT3-MoE by substituting the MLP with MoE in the second Transformer block of every consecutive pair, while implement LLaMA2-MoE by by substituting the MLP with MoE in every Transformer block. For models undergoing zero-shot evaluation on downstream tasks such as HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), BoolQ (Clark et al., 2019), ARC-Easy (Clark et al., 2018), OpenBookQA (Mihaylov et al., 2018), RACE (Lai et al., 2017), and MathQA (Amini et al., 2019), we pre-train the models using various architectures on a 1B token subset of the SlimPajama-627B dataset (Soboleva et al., 2023). For models evaluated on WikiText-103 (Merity et al., 2017), we conduct pre-training with different architectures on the OpenWebtext dataset (Gokaslan & Cohen, 2019). Table 8 summarizes the hyperparameters for training the GPT2-MoE-Small, GPT2-MoE-Medium, GPT3-MoE-XL and LLaMA-MoE models.
A.9 Additional Examples of Intermediate Representations Similarities
Figure 14: Intermediate representation similarities in LLaMA2-MoE.
Figure 15: Intermediate representation similarities in OLMoE (Muennighoff et al., 2024).